# Test-time Adaptive Vision-and-Language Navigation

### v2

@@ Abstract
Vision-and-Language Navigation (VLN) has witnessed significant advancements in recent years, largely attributed to meticulously curated datasets and proficiently trained models. Nevertheless, when tested in diverse environments, the trained models inevitably encounter significant shifts in data distribution, highlighting that relying solely on pre-trained and fixed navigation models is insufficient. To enhance models’ generalization ability, test-time adaptation (TTA) demonstrates significant potential in the computer vision field by leveraging unlabeled test samples for model updates. However, simply applying existing TTA methods to the VLN task cannot well handle the adaptability-stability dilemma of VLN models, i.e., frequent updates can result in drastic changes in model parameters, while occasional updates can make the models ill-equipped to handle dynamically changing environments. Therefore, we propose a Fast-Slow Test-Time Adaptation (FSTTA) approach for VLN by performing decomposition-accumulation analysis for both gradients and parameters in a unified framework. Specifically, in the fast update phase, gradients generated during the recent multi-step navigation process are decomposed into components with varying levels of consistency. Then, these components are adaptively accumulated to pinpoint a concordant direction for fast model adaptation. In the slow update phase, historically recorded parameters are gathered, and a similar decomposition-accumulation analysis is conducted to revert the model to a stable state. Extensive experiments show that our method obtains impressive performance gains on four popular benchmarks.
@@ 1 Introduction
Developing intelligent agents capable of adhering to human directives remains a significant challenge in embodied AI. Recently, Vision-and-Language Navigation (VLN) [3, 52, 11, 38, 54, 79], which requires an agent to comprehend natural language instructions and subsequently execute proper actions to navigate to the target location, serves as a useful platform for examining the instruction-following ability. Despite tremendous progress has been achieved such as transformer-based sequence-to-sequence learning [25, 9, 11], large-scale training data collection [10, 77], and various reinforcement and imitation learning strategies [74, 17], the navigational capabilities of agents within varied testing environments still warrant further improvement.
@@
Figure 1: (a) Illustration of the data distribution shift between training and testing samples in the VLN Task. (b) Comparison between various TTA strategies on REVERIE [52] validation unseen set using OSR and SR metrics. Among them, ‘DUET’ [11] is the base model, ‘Frequent Update’ means updating at certain intervals within each sample, ‘Stable Update’ refers to initializing with the original base model for each sample and using its best in-sample update interval INT=1. All these strategies adopt TENT [68] for model updates. The results show that overly fast or overly slow TTA fail to achieve significant improvements.
In the VLN task, agents are required to sequentially execute actions contingent upon the evolving environmental cues. Regrettably, owing to disparities in environmental factors, such as distinct room types and objects as shown in Figure 1(a), the trained agents inevitably confront significant shifts in data distribution when applied in practical scenarios [20, 22]. In light of this issue, depending solely on a pre-trained and fixed VLN model is inadequate.
@@
Recently, Test-Time Adaptation (TTA) [39, 40, 51, 71] has been recognized as an effective technique for leveraging unlabeled test samples to update models and address shifts in data distribution. It has garnered notable success across various computer vision tasks, such as image classification [68, 51], segmentation [71, 67], and video classification [42, 80]. For instance, TENT [68] utilizes an entropy minimization objective to update model parameters, thereby enhancing the generalization ability to recognizing test data. Nonetheless, the application of TTA in the realm of VLN remains relatively uncharted. Although prevailing TTA methodologies can be integrated into VLN models with certain alterations, this direct application cannot well handle the adaptability-stability dilemma of models due to the multi-step action-execution nature of VLN. Specifically, in contrast to traditional classification tasks, where a single TTA operation suffices for a test sample, VLN mandates an agent to perform sequential actions within a single test sample. On one hand, while conducting TTA at every (or a few) action steps enables rapid agent adaptation to dynamic environments, frequent model updates may introduce significant model alterations, potentially causing cumulative errors and catastrophic forgetting [71, 50, 60], thus compromising model stability during testing. On the other hand, initializing the same model for stable TTA in each test sample may hinder the model’s ability to adaptively learn experience from historical test samples, thereby impeding its potential for achieving superior performance.  Figure 1(b) shows that both overly fast or overly slow model updates fail to achieve significant performance improvements.
@@
To tackle the above issues, we proposes a Fast-Slow Test-Time Adaptation (FSTTA) method for the VLN tasks. Built upon a unified gradient-parameter decomposition-accumulation framework, our approach consists of a fast update phase and a slow update phase, pursuing a balance between adaptability and stability in model updates. Specifically, with a test-time training objective, such as entropy minimization, we can derive gradients at each action step in the fast update phase. However, due to the unsupervised nature of TTA, these gradients inevitably contain noise information. Using these gradients for model update can interfere with the adaptability, especially when the update is frequently invoked. Therefore, we attempt to find a reliable optimization direction by periodically analyzing the gradients generated during the recent multi-step navigation process. We first establish a local coordinate system to decompose these gradients into components with varying levels of consistency. Subsequently, these components are adaptively accumulated to pinpoint a concordant direction for updating the model. Besides, a gradient variance regularization is incorporated to dynamically adjust the learning rate.
After a certain number of fast updates, the model parameters (also called model state) are recorded. To further mitigate the issues of cumulative errors and catastrophic forgetting that may result from excessively frequent model updates, during the slow update phase, we revert the model to its historical state and conduct a decomposition-accumulation analysis on the parameter variation trajectory for a direct model update. This process is akin to the fast phase but shifts its focus from gradients to the parameters. Both phases are performed alternately during testing to balance the adaptability and stability of the model. As shown in Figure 1(b), the proposed method achieves significant improvement against other model update strategies.
Our contributions can be summarized as follows:
  • $\bullet$
    We investigate the test-time adaptation within the realm of VLN. Our proposed method verifies TTA as a promising and viable avenue for enhancing VLN performance.
  • $\bullet$
    Based on a unified decomposition-accumulation framework for both gradients and parameters, our method ensures swift model adaptability to environmental changes in the short-term fast update phase, while preserves stability throughout the long-term slow update phase.
  • $\bullet$
    Our FSTTA elevates the performance of several leading VLN models across four popular benchmarks. When applied to the notable DUET model [11], our method yields a performance boost of over $5\%$ on the representative discrete/continuous datasets REVERIE/R2R-CE. Furthermore, our method shows superior results compared to other premier TTA techniques.
@@ 2 Related Work
Vision-and-Language Navigation (VLN). Recently, VLN tasks have received significant attention in the embodied AI domain and a variety of effective approaches have been proposed [3, 86, 20]. Most existing methods facilitate VLN research by developing powerful techniques for model training, including: (i) Designing advanced network architectures. Sequence-to-sequence framework are the most commonly used one for predicting agent actions from a historical observation series. Early VLN models utilize LSTMs with various attention mechanisms [3, 16, 46, 24] while recent ones [23, 25, 9, 41, 11, 27, 1] resort to the more popular transformer-based methods for multi-modal pre-training. Other architectures are also explored such as graph neural networks [86] and parameter-efficient adapter [55]. (ii) Adopting various training paradigms such as reinforcement and imitation learning [49, 63, 74, 17]. Moreover, to estimate the completeness of instruction following and decide when to conduct backtracking, progress monitoring [45, 85] and back-tracking [30, 46] are also employed to promote training process. (iii) Performing data augmentation for training a stronger model. In recent years, more and more large-scale benchmarks are established via collecting human annotations [33, 86, 57] or creating new environments [52, 10]. Other approaches explore techniques such as mixup and synthesis [43, 29], style transfer [37], or future-view image semantics [36] for data augmentation. (iv) Leveraging additional information for boosting model capacity. Since the goal of VLN is to navigate in photo-realistic environments, there are many kinds of information in the world that can be used such as knowledge [38], 3D scene geometry [44, 78], and landmarks [72, 12]. Though the above approaches have attempted to employ fruitful strategies for training an effective VLN model, they still struggle to adequately address the domain discrepancy between training and testing data. Some studies [55, 54] have tried to use dynamic networks or auxiliary models to cope with varying test data, however, these approaches neither directly minimize the domain gap between the model and test data nor offer dynamic model update strategies.
@@
Test-time Adaptation (TTA). TTA, which allows models to adapt the test data in an online and unsupervised manner, has gained a lot of interest with a host of various approaches proposed in the literature [39, 40, 61, 73, 34]. Existing TTA methods generally rely on batch normalization calibration [48, 84, 19], entropy minimization [68, 51, 50, 64], auxiliary self-supervised task or data regularization [62, 28, 5, 83, 66] to acquire useful information for reducing the domain gap between training and testing data. To stabilize adaptation in continuously changing data distribution, recently, continual test-time adaptation [71, 50, 60, 6, 13, 82], as a more practical setting, has been tentatively explored for addressing the cumulative errors and catastrophic forgetting issues. Until now, test-time adaptation has been preliminarily explored in some sequential data analysis fields such as action recognition [42] and video classification [80]. However, TTA on VLN tasks is yet to be explored.
@@
Gradient-based Methods. Gradients are typically central to modern SGD-based deep learning algorithms. To date, gradient analysis research has predominantly focused on domain generalization (DG) [47, 35, 70, 56, 76, 65], due to the negative impact of conflicting gradients from multiple domains on model optimization. Pioneering works [15, 81, 47] perform gradient surgery at the backpropagation phase via various strategies such as normal plane projection [81] and consensus learning [47]. Other approaches resort to gradient agreement regularization for refining the optimization direction by leveraging sharpness [70] or similarity [58, 56] measurements. Different from the above models that only consider a single-phase gradient surgery in DG, we jointly analyze the gradient-parameter states for a two-phase (fast-slow) TTA in the VLN task.
@@ 3 Our Approach
Figure 2: Overall framework of the proposed Fast-Slow Test-Time Adaptation (FSTTA) for VLN tasks. In the fast update phase, taking ‘Sample i’ as an example, the model periodically analyzes the gradients ($\{\bm{g}\}$) generated during the recent multi-step navigation and performs a gradient decomposition-accumulation analysis to pinpoint a concordant direction for model update. After a certain number of fast updates, historical model parameters ($\{\bm{\Theta}\}$) are recorded. In the slow update phase, we revert the model to its historical state and conduct a parameter decomposition-accumulation analysis to learn an optimization path for direct parameter modulation. Note that ‘F’, ‘S’ in the robots means the model parameters after fast and slow updates. ‘$\text{F}_{1}$’ indicates the first fast update within a test sample.
3.1 Preliminaries and Framework Overview
Problem Setup and VLN Base Model. Given a natural language instruction $\bm{I}$, the VLN task requires an agent to find the target viewpoint through the environment by executing a series of actions. During the navigation process, an undirected exploration graph $\mathcal{G}_{t}=(\mathcal{V}_{t},\mathcal{E}_{t})$ is progressively constructed, where $\mathcal{V}_{t}$ denotes navigable nodes, $\mathcal{E}_{t}$ indicates the connectivity edges, $t$ is the current timestep. At this moment, the agent receives a panoramic view that contains 36 single images. The panorama is represented by the image features $\mathcal{R}_{t}$ and their object features $\mathcal{O}_{t}$, where these features can be extracted by pre-trained vision transformers (ViT) [14, 11, 38]. To accomplish the instruction, the agent needs to predict probabilities for the currently navigable nodes and select the most possible one as the next movement action. The probabilities can be predicted as:
@@ 3.1 Preliminaries and Framework Overview
where $\mathcal{H}_{t}$ indicates the history information that encodes the observed visual features and performed actions [11, 55]. $\phi(\cdot)$ is the VLN base model such as the dual-scale graph transformer [11, 10], $\bm{\Theta}$ is the learnable model parameters.
Framework Overview. In this paper, we devote to adjusting the VLN base model during testing process within an unsupervised manner. Our FSTTA framework is illustrated in  Figure 2. For each sample, at timestep $t$, we employ the commonly adopted entropy minimization objective [68, 51] for test-time adaption, which aims to reduce the entropy of the probabilities over the current navigable nodes:
@@
During the optimization of the above objective, gradients are back-propagated for updating the model’s parameters. However, updating the whole base model is computationally infeasible. As a result, we only consider a small portion of the model parameters for gradient calculation. Since affine parameters in normalization layers capturing data distribution information, numerous TTA methods opt to update these parameters for adaption [68, 51, 39]. In this paper, we employ the model’s final a few layer-norm operations for TTA and maintain other parameters frozen. For brevity, we still use the symbol $\bm{\Theta}$ to represent these parameters to be updated, $\bm{\Theta}\in\mathbb{R}^{D}$. Targeting at fully leveraging the gradient and parameter information, under a unified decomposition-accumulation analysis framework, we propose an effective two-phase adaptation for fast and slow model updates.

### v3

# Fast-Slow Test-Time Adaptation for Online Vision-and-Language Navigation

@@ Abstract
The ability to accurately comprehend natural language instructions and navigate to the target location is essential for an embodied agent. Such agents are typically required to execute user instructions in an online manner, leading us to explore the use of unlabeled test samples for effective online model adaptation. However, for online Vision-and-Language Navigation (VLN), due to the intrinsic nature of inter-sample online instruction execution and intra-sample multi-step action decision, frequent updates can result in drastic changes in model parameters, while occasional updates can make the model ill-equipped to handle dynamically changing environments. Therefore, we propose a Fast-Slow Test-Time Adaptation (FSTTA) approach for online VLN by performing joint decomposition-accumulation analysis for both gradients and parameters in a unified framework. Extensive experiments show that our method obtains impressive performance gains on four popular benchmarks. Code is available at https://github.com/Feliciaxyao/ICML2024-FSTTA.
Machine Learning, ICML
@@ 1 Introduction
Developing intelligent agents capable of adhering to human directives remains a significant challenge in embodied AI. Recently, Vision-and-Language Navigation (VLN) (Anderson et al., 2018; Qi et al., 2020; Chen et al., 2022d; Yang et al., 2023), which requires an agent to comprehend natural language instructions and subsequently perform proper actions to navigate to the target location, serves as a useful platform for examining the instruction-following ability.
@@
Figure 1: (a) Illustration of online VLN. (b) Comparison between TTA strategies on REVERIE (Qi et al., 2020) validation unseen set using SPL and SR metrics. ‘DUET’ (Chen et al., 2022d) is the base model, ‘Frequent Update’ means updating at certain intervals within each sample, ‘Stable Update’ refers to initializing with the original base model for each sample and using its best intra-sample update interval INT=1. All these strategies adopt TENT (Wang et al., 2021) for model updates. The results show that overly fast or overly slow TTA fail to achieve significant improvements.
In practical applications, a trained VLN agent is required to execute user instructions in various environments at different times in an online manner, as depicted in Figure 1(a). However, owing to disparities in environmental factors, such as distinct room types and objects, the trained agent inevitably confront data discrepancies during online testing (Gu et al., 2022; Guhur et al., 2021). This raises an important question: can an agent accumulate experience and enhance its capabilities while executing instructions? However, due to the lack of annotation, updating the model via supervised online learning is impractical. In addition, other learning paradigms like unsupervised domain adaptation or semi-supervised learning are also infeasible, given considerations for the issues of execution efficiency and privacy protection.
@@
Recently, online Test-Time Adaptation (TTA) (Liang et al., 2023; Niu et al., 2023; Wang et al., 2022a) has been recognized as an effective technique of online model updating by leveraging unlabeled test samples. Although prevailing TTA methods can be integrated into VLN models with certain alterations, this direct application cannot well handle the adaptability-stability dilemma of models due to the intrinsic nature of inter-sample online instruction execution and intra-sample multi-step action decision. Specifically, unlike traditional classification tasks where a single TTA operation is sufficient for each test sample, as illustrated in Figure 1(a), online VLN requires an agent to perform a sequence of actions within a single test sample and to execute various samples (instructions) in an online manner. On one hand, while conducting TTA at every (or a few) action steps enables rapid agent adaptation to dynamic environments, frequent model updates may introduce significant model alterations, potentially causing cumulative errors and catastrophic forgetting (Wang et al., 2022a; Niu et al., 2022; Song et al., 2023), thus compromising model stability during testing. On the other hand, initializing the same model for stable TTA in each test sample may hinder the model’s ability to adaptively learn experience from historical test samples, thereby impeding its potential for achieving superior performance. Figure 1(b) shows that both overly fast or overly slow model updates fail to achieve significant performance improvements.
@@
To tackle the above issues, we propose a Fast-Slow Test-Time Adaptation (FSTTA) method for online VLN tasks. Built upon a unified gradient-parameter decomposition-accumulation framework, our approach consists of a fast update phase and a slow update phase, pursuing a balance between adaptability and stability in online model updating. Specifically, with a test-time training objective, such as entropy minimization (Wang et al., 2021), we can derive gradients at each action step in the fast update phase. However, due to the unsupervised nature of TTA, these gradients inevitably contain noise information. Using these gradients for model update can interfere with the adaptability, especially when the update is frequently invoked. Therefore, we establish a local coordinate system to find a reliable optimization direction by periodically analyzing the gradients generated during the recent multi-step navigation process. After a certain number of fast updates, the model parameters (also called model state) are recorded. To further mitigate the issues of cumulative errors and catastrophic forgetting that may result from excessively frequent model updates, during the slow update phase, we revert the model to its historical state and conduct a decomposition-accumulation analysis on the parameter variation trajectory for a direct model update. Both phases are performed alternately during testing to balance the adaptability and stability of the model. As shown in Figure 1(b), the proposed method achieves significant improvement against other model update strategies.
Our contributions can be summarized as follows: (1) Considering the characteristics of inter-sample online instruction-execution and intra-sample multi-step action-execution, we explore the online VLN task and propose a fast-slow test-time adaptation (FSTTA) method for effective navigation and model updating. (2) Based on a unified decomposition-accumulation framework for both gradients and parameters, our method ensures swift model adaptability to environmental changes in the short-term fast update phase, while preserves stability throughout the long-term slow update phase. (3) Our FSTTA elevates the performance of several leading VLN models on four popular benchmarks. When applied to the notable DUET model (Chen et al., 2022d), it yields a performance boost of over $5\%$ on the representative discrete/continuous benchmarks REVERIE/R2R-CE. Furthermore, our method shows superior results compared to other premier TTA techniques.
@@ 2 Related Work
Vision-and-Language Navigation (VLN). Most existing methods facilitate VLN research by developing powerful techniques for model training, including: (i) Designing advanced network architectures. Early VLN models utilize LSTMs with various attention mechanisms (Anderson et al., 2018; Fried et al., 2018; Hong et al., 2020) while recent ones (Hao et al., 2020; Hong et al., 2021; Chen et al., 2021; Lin et al., 2022; Chen et al., 2022d; Huo et al., 2023) resort to the more popular transformer-based methods for multi-modal pre-training. Other architectures are also explored such as graph neural networks (Zhu et al., 2021) and parameter-efficient adapter (Qiao et al., 2023b). (ii) Adopting various training paradigms such as reinforcement and imitation learning (Nguyen et al., 2019; Tan et al., 2019; Wang et al., 2019; Gao et al., 2023a). Moreover, to estimate the completeness of instruction following and decide when to conduct backtracking, progress monitoring (Ma et al., 2019a; Zhu et al., 2020) and back-tracking (Ke et al., 2019; Ma et al., 2019b) are also employed to promote training process. (iii) Performing data augmentation for training a stronger model. In recent years, more and more large-scale benchmarks are established via collecting human annotations (Ku et al., 2020; Zhu et al., 2021; Ramrakhya et al., 2022) or creating new environments (Qi et al., 2020; Chen et al., 2022c). Other approaches explore techniques such as mixup and synthesis (Liu et al., 2021; Kamath et al., 2023), style transfer (Li et al., 2022), or future-view image semantics (Li & Bansal, 2023) for data augmentation. (iv) Leveraging additional information for boosting model capacity. Since the goal of VLN is to navigate in photo-realistic environments, there are many kinds of information in the world that can be used such as knowledge (Li et al., 2023), 3D scene geometry (Liu et al., 2023; Wang et al., 2023f), and landmarks (Wang et al., 2022b; Cui et al., 2023).
Although the above methods have made significant progress in training effective models, they overlook the utilization of test data during the online VLN process. In real-world applications, an agent is required to continuously execute user instructions at different times. The ability of an agent to accumulate experience during this process would greatly enhance its practical value. Note that (Lu et al., 2022) first explored test-time adaptation for VLN. However, this approach does not perform model update in an online manner and overlook the balance between adaptability-stability.
@@
Online Test-time Adaptation (TTA). TTA allows models to adapt the test data in an online and unsupervised manner (Liang et al., 2023; Lim et al., 2023; Lee et al., 2023). Existing TTA methods generally rely on batch normalization calibration (Mirza et al., 2022; Zhao et al., 2023; Gong et al., 2022), entropy minimization (Wang et al., 2021; Niu et al., 2023; Tang et al., 2023), auxiliary self-supervised task or data regularization (Sun et al., 2020; Boudiaf et al., 2022; Zhang et al., 2022) to acquire useful information for reducing the domain gap between training and testing data. To stabilize adaptation in continuously changing data distribution, recently, continual test-time adaptation (Wang et al., 2022a; Niu et al., 2022; Song et al., 2023; Döbler et al., 2023; Yuan et al., 2023; Liu et al., 2024), as a more practical setting, has been tentatively explored for addressing the cumulative errors and catastrophic forgetting issues. Until now, test-time adaptation has been preliminarily explored in some sequential data analysis fields such as action recognition (Lin et al., 2023) and video classification (Yi et al., 2023). However, they overlook the joint inter- and intra-sample structure in TTA of sequential data.
@@
Gradient-based Methods. Gradients are central to modern SGD-based deep learning algorithms. To date, gradient analysis research has predominantly focused on domain generalization (DG) (Mansilla et al., 2021; Lew et al., 2023; Wang et al., 2023b; Rame et al., 2022; Wang et al., 2023e; Tian et al., 2023), due to the negative impact of conflicting gradients from multiple domains on model optimization. Pioneering works (Du et al., 2018; Yu et al., 2020; Mansilla et al., 2021) perform gradient surgery at the backpropagation phase via various strategies such as normal plane projection (Yu et al., 2020) and consensus learning (Mansilla et al., 2021). Other approaches resort to gradient agreement regularization for refining the optimization direction by leveraging sharpness (Wang et al., 2023b) or similarity (Shi et al., 2021; Rame et al., 2022) measurements. Different from the above strategies that only consider a single-phase gradient surgery in DG, we jointly analyze the gradient-parameter states for a two-phase (fast-slow) TTA in the VLN task.
@@ 3 Our Approach
Figure 2: Overall framework of the proposed Fast-Slow Test-Time Adaptation (FSTTA) for online VLN. In the fast update phase, taking ‘Sample i’ as an example, the model periodically analyzes the gradients ($\{\bm{g}\}$) generated during the recent multi-step navigation and performs a gradient decomposition-accumulation analysis to pinpoint a concordant direction for model update. After a certain number of fast updates, historical model parameters ($\{\bm{\Theta}\}$) are recorded. In the slow update phase, we revert the model to its historical state and conduct a parameter decomposition-accumulation analysis to learn an optimization path for direct parameter modulation. Note that ‘F’, ‘S’ in the robots means the model parameters after fast and slow updates. ‘$\text{F}_{1}$’ indicates the first fast update within a test sample.
Problem Setup and VLN Base Model. Given a natural language instruction $\bm{I}$, the VLN task requires an agent to find the target viewpoint through the environment by executing a series of actions. During the navigation process, an undirected exploration graph $\mathcal{G}_{t}=(\mathcal{V}_{t},\mathcal{E}_{t})$ is progressively constructed, where $\mathcal{V}_{t}$ denotes navigable nodes, $\mathcal{E}_{t}$ indicates the connectivity edges, $t$ is the current timestep. Note that a ‘STOP’ node is added to this graph to indicate a stop action, and we connect it with all other nodes. At this moment, the agent receives a panoramic view that contains 36 single images. The panorama is represented by the image features $\mathcal{R}_{t}$ and their object features $\mathcal{O}_{t}$, where these features can be extracted by pre-trained vision transformers (ViT) (Dosovitskiy et al., 2020; Chen et al., 2022d; Li et al., 2023). To accomplish the instruction, the agent needs to predict probabilities for the currently navigable nodes and select the most possible one as the next movement action. The probabilities can be predicted as:
@@
where $\mathcal{H}_{t}$ indicates the history information that encodes the observed visual features and performed actions (Chen et al., 2022d; Qiao et al., 2023b). $\phi(\cdot)$ is the VLN base model such as the dual-scale graph transformer (Chen et al., 2022d, c), $\bm{\Theta}$ is the learnable model parameters.
Framework Overview. In this paper, we devote to adjusting the VLN base model during testing process within an online and unsupervised manner. Our FSTTA framework is illustrated in  Figure 2. For each sample, at timestep $t$, we employ the commonly adopted entropy minimization objective (Wang et al., 2021; Niu et al., 2023) for test-time adaption, which aims to reduce the entropy of the probabilities over the current navigable nodes:
@@
During the optimization of the above objective, gradients are back-propagated for updating the model’s parameters. However, updating the whole base model is computationally infeasible. As a result, we only consider a small portion of the model parameters for gradient calculation. Since affine parameters in normalization layers capturing data distribution information, numerous TTA methods opt to update these parameters for adaption (Wang et al., 2021; Niu et al., 2023; Liang et al., 2023). In this paper, we employ the model’s a few final layer-norm operations for TTA and maintain other parameters frozen. For brevity, we still use the symbol $\bm{\Theta}$ to represent these parameters to be updated in the following sections, $\bm{\Theta}\in\mathbb{R}^{D}$. Targeting at fully leveraging the gradient and parameter information, under a unified decomposition-accumulation analysis framework, we propose an effective two-phase adaptation for fast and slow model updates.